1- if I have angle A = angle B and angle D = angle C
Prove that lines AB parallel to line DC.
1
Expert's answer
2010-12-11T18:28:10-0500
A=B and C=D, and A+B+C+D=360
2A + 2 C = 360;
A+C = 180.
The sum of unilateral internal angles equals to 180°, then the lines are parallel.